OpenStudy (anonymous):

a man bought a tv that was listed at 160$.he was given successive discount of 20% and 10%.The price he paid was?

OpenStudy (anonymous):

160*.20=32 160-32=128 128*.10=12.8 128-12.8=107.2 107.2

OpenStudy (anonymous):

the final answer is 105.2 and not 107.2 but the solution is correct.

OpenStudy (anonymous):

*128-12.8=115.2 sorry. and not 105.2